OpenStudy (anonymous):

Differentiate the expression x^5y^5 with respect to x. (Use D for dy/dx). Why do I not get this would it not be (5x^4)/(5y^4)

OpenStudy (anonymous):

you need the product rule for this because it is a product

OpenStudy (anonymous):

f1+g+f+g1 right

OpenStudy (anonymous):

\[(fg)'=f'g+g'f\] with \[f(x)=x^5, f'(x)=5x^4, g(x)=y, g'(x)=y'\]

OpenStudy (anonymous):

(5x^4)(y^5)+(x^5)(5y^4)

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

\[5x^4y^5+5x^5y^4y'=0\] solve for \(y'\)

OpenStudy (anonymous):

so subtract the y over and turn everything negative?

OpenStudy (anonymous):

you subtract \(5x^4y^5\) from both sides, then divide both sides by \(5x^5y^4\)

OpenStudy (anonymous):

so -5x^5y^4/5x^5y^4

OpenStudy (anonymous):

If it is given x^5y^5=0,then \[5x ^{4}y ^{5}+x ^{5}*5y ^{4} D=0\] \[5x ^{5}y ^{4}D=-5x ^{4}y ^{5}\] \[D=-\frac{ 5x ^{4}y ^{5} }{ 5x ^{5}y ^{4} }=-\frac{ y }{ x }\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw